Now we will calculate the actual probability  that all the representatives that chosen for presentations will be women. We have found that the probability is given as

P=(10C8/15C8)


Express your answer as a fraction in simplest form.
Answer #1

from given statement : (10C8) =10!/(8!*(10-8)!) =10*9/(2*1) =45

(15C8) =15!/(8!*(15-8)!) =(15*14*13*12*11*10*9)/(7*6*5*4*3*2*1)=6435

therefore P=(10C8/15C8) =45/6435 =1/143